OpenStudy (anonymous):

Write the explicit formula for the sequence with a1 = –4 and r = –7.

OpenStudy (anonymous):

SO that is in which series??

OpenStudy (anonymous):

a_n = 90 • 4^(n - 1) for n = 1, 2, ... 90, 360, 1440, ...

OpenStudy (anonymous):

A. an = –4 • (–7)n – 1 B. an = –4 • 7–n – 1 C. an = –(4 • 7)n – 1 D. none of these

OpenStudy (anonymous):

D

OpenStudy (anonymous):

when r is given then the sequence is Geometric Progression because r represents common ratio so general term in GP would be \[t _{n} = a1*r^{n-1}\]